OpenStudy (anonymous):

We are interested in the dimensions of a certain square. A rectangle has length twice the side of this square and width three units less than the side of this square. Which equation describes this situation if the two areas are equal?

OpenStudy (anonymous):

does it say the length width and height of the original

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok I'll see what I can do

OpenStudy (anonymous):

its asking the step you would use

OpenStudy (anonymous):

ok I see so r there any options or is it essay question

OpenStudy (anonymous):

x^2=2x(x-3) x^2=2x^2-6 x^2-2x+6=0 2x^2-x+6=0

OpenStudy (anonymous):

the second to last one C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem it helps to help other people

OpenStudy (anonymous):

We are interested in the dimensions of a certain rectangle. This rectangle has length 5 units more than the side of this square and width half the side of this square. If the two areas are equal, what are the rectangle's dimensions (w x h)?

OpenStudy (anonymous):

r thee any choices sorry it help me so I can fill them in

OpenStudy (anonymous):

6 x 9/2 12 x 3/2 10 x 5/2 8 x 7/2

OpenStudy (anonymous):

thank you the answer is C again lol

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem